OpenStudy (anonymous):

x squared + 6x+11=0 completing the square show steps.

OpenStudy (cwrw238):

x^2 + 6x + 11 = 0 (x + 3)^2 - 9 + 11 = 0 (x + 3)^2 = -2 now take square root of both sides to solve